# Enhancing Training Data Attribution with Representational Optimization

### v1

@@ 5 Evaluation of Data Attribution
To evaluate these data attribution methods, we follow [1] and use the linear datamodeling score (LDS) as our evaluation metric. For each evaluation dataset, we sample 100 different random subsets $(S_{1},\cdots,S_{100})$ of its training set, each containing $n$ training examples, and train the target LM on each of these subsets. For each example $x$ in the test set, we approximate the expectation of the model prediction $r(x,S_{i})$ as the evaluation cross-entropy loss. Given the score $f(x,S_{i})$ predicted by the data attribution model, the LDS is computed as the Spearman rank correlation between the true and estimated influence scores: $\rho\bigl{(}\{r(x,S_{i}):i\in[100]\},\{f(x,S_{i}):i\in[100]\}\bigr{)}$. Finally, the LDS on the dataset is averaged across all test examples.
@@ Results
Table 1 presents the LDS evaluation results for attributions computed on the Qwen2.5-0.5B model’s predictions across four evaluation datasets. AirRep achieves the best performance across all evaluation datasets. Notably, it outperforms the competitive gradient-based approach LoGra [16], despite being 80$\times$ more computationally efficient and 50$\times$ more storage efficient. Additionally, previous representation-based methods exhibit low correlation scores in this evaluation, underscoring the importance of task-aware optimization of representations.
Figure 3 illustrates the LDS correlation when applying TDA methods to different target language models, ranging from Qwen2.5-0.5B to Qwen2.5-7B. For gradient-based methods such as LoGra and LESS, the proxy model can be adjusted—i.e., a smaller LM can be used as a proxy to compute influence scores for larger target LMs [25]. Additionally, the projection dimensionality can be modified to balance computation, storage, and performance. In all cases, we use the same AirRep model trained exclusively on data generated by Qwen2.5-0.5B. Notably, AirRep demonstrates strong performance when applied to larger target LMs. Specifically, it consistently achieves results comparable to the best gradient-based methods, which require significantly more resources (e.g., 84$\times$ to 262$\times$ more computation time and 50$\times$ to 80$\times$ more storage). These results highlight AirRep’s potential for efficient and effective data attribution. AirRep also demonstrates strong generalization across different data sizes and target LM types (Figure 7 and Figure 8).
@@ 7 Evaluation of Data Classification
TDA methods have been applied to measuring data similarity for the purpose of model explainability [6]. To evaluate the effectiveness of such explanations, Hanawa et al. [6] introduced the “Identical Class Test” for image classification, which assesses whether the most influential training example identified by TDA methods belongs to the same class as the test instance.
@@ Results
Table 2 presents the evaluation results of LoGra, GTE-Small, and AirRep. The results indicate that AirRep significantly improves data classification accuracy on FLAN compared to GTE-Small (increasing from 50.59 to 86.41) and also outperforms LoGra. Notably, since AirRep is trained without using any data labels, this suggests that the model can learn to represent task-related information in an unsupervised manner. A similar trend is observed for Tulu, where AirRep demonstrates superior performance in identifying the data source. For SafeRLHF, AirRep outperforms LoGra but lags behind GTE-Small, likely because AirRep’s training data, generated from UltraChat, does not contain harmful content and thus lacks the necessary learning signal.
@@ 9 Amortizing Training Cost
Compared with gradient-based methods, AirRep is more efficient at inference but requires training on auto-generated data, which incurs additional cost. Our evaluation above shows that FLAN- or UltraChat-pretrained AirRep generalizes well on new TDA data and tasks. We now consider another setting where AirRep is re-trained from scratch for each TDA task, and analyze how its inference efficiency can amortize its training cost [17].
@@
Then, Figure 6 (right) compares the number of training examples that can be processed under different A100 GPU hour budgets by AirRep (crossover checkpoint) and LoGra, accounting for both data generation and model training costs. LoGra completes model training earlier, but AirRep—despite longer training time—achieves nearly two orders of magnitude more throughput once model is trained. Notably, a turning point appears around 475K examples, where both methods take 5.6 GPU hours. Beyond this, even with retraining, AirRep is faster than LoGra. For example, with 24 GPU hours, AirRep can attribute over 100M examples, whereas LoGra can process only around 2M. This highlights the advantage of AirRep for large-scale TDA scenarios. We additionally emphasize that, as shown in Section 5, AirRep can be transferred from a small model to larger models and still attain stronger performance than all baselines, so the “re-training from scratch” results in this section represent worst-case results for AirRep.
@@ 10 Related Work
Most existing work in TDA focuses on gradient-based methods, such as Influence Functions [3] and their subsequent optimizations, including group influence estimation [41, 13], Hessian approximation [42, 21, 43], gradient normalization [24, 25], gradient projection [22, 16], model ensembling [12, 33, 44], and distillation of the proxy influence models [45, 46, 47]. However, studies have found that the derivation of gradient-based methods may rely on incorrect assumptions in large neural networks [48, 4]. Additionally, gradient-based methods suffer from the computational expense of gradient calculations and Hessian inverses [21, 33].
Alternatively, representation-based approaches for TDA measure data influence in the representation space [7, 49, 6, 9], offering computational and storage efficiency. However, current methods rely on heuristically designed representations without task-aware and model-specific optimization, often leading to suboptimal results [12]. Some research on TDA explores simulation-based methods such as datamodels  [50, 51, 52, 1], which are usually not tractable for LLM applications due to their computational expense.
TDA methods have been applied to various purposes in machine learning, such as identifying mislabeled training data points [3, 53], detecting data poisoning attacks [54, 55], and guiding data augmentation [56, 57]. TDA methods have also been used for understanding LLM generalization [21] and learning processes [30, 58]. Another notable line of related work is data selection for LLMs, which focuses on curating data to optimize test performance [59]. TDA-like functions for measuring relationships between data have been widely adopted in data selection pipelines to enhance the quality and diversity of the selected data. Both gradient-based [60, 25, 47] and representation-based [8, 61, 62, 10] TDA approaches have been applied. Meanwhile, studies have explored using text embedding models to select in-context examples for LLM prompting [63, 64, 65].
@@ 11 Conclusion
This paper introduces a new approach to training data attribution called AirRep, which combines the advantages of task-driven optimization in gradient-based approaches with the efficiency of representation-based methods. AirRep represents training data as embeddings and scores their influence using inner-product similarity in the representation space. AirRep’s novel attention-based pooling mechanism captures the group effects, and the entire model is optimized using a novel proposed weighted pairwise ranking objective over automatically generated influence signals. Evaluation on LLM fine-tuning demonstrates the effectiveness and generalization of AirRep. Limitations of this work include the added cost of model training and evaluation limited to the LLM fine-tuning stage. For future work, we plan to extend our method to additional scenarios, such as the pre-training stages of LLMs and multi-modal model training.
@@ B.1 Group Influence Function
Extended from Basu et al. [13] results, we show that the structure of the high-order group influence function is given by:
@@ B.2 Deriving the First-Order Group Influence Function
[…]
@@ B.4 Deriving the Third-Order Group Influence Function
For many practical purposes, second-order corrections already capture the leading “beyond first-order” effects. However, in order to have a more complete perception of the structure of the influence function (and also to address our curiosity), we will continue to derive the third-order group influence function.
@@
[…]
@@
[…]
  • Term (iii) vanishes because the per-sample weight is linear in $\epsilon$:
[…]
@@
[…]
@@
  • Term (v) involves the third derivative with respect to $\theta$. In most influence-function derivations, one often assumes that higher-order derivatives $\nabla^{3}_{\theta}L(\theta^{*})$ are negligible. Following this convention, we consider:
[…]
@@
[…]
@@ B.5 Deriving the Higher-Order Group Influence Function
We see that at each order $k$ in $\epsilon$, the only “new” term that enters the equations (and thus the recursion for $\Delta_{k}$) is the first derivative with respect to $\epsilon$ multiplied by lower-order $\Delta_{j}$’s, plus the usual expansions in powers of $\Delta_{\epsilon}$ (where $\Delta_{\epsilon}=\theta_{\epsilon,S}^{*}-\theta^{*}$).
@@ Appendix C Comparison between gradient-based and representation-based methods
  • Computation cost: Gradient-based methods generally rely on the computation of gradients and some approximation of the Hessian, making them computationally expensive. In practice, the cost of gradient-based methods is comparable to training the model on the all of the target examples because computing influence involves computing loss gradients for each target example. In contrast, representation-based methods are efficient and applicable for large-scale retrieval [66].
  • Storage cost: Influence functions can be particularly storage-intensive as they can involve storing the full gradient (equal in size to the model parameters themselves) for each data sample, which prevents large-scale data attribution. Follow-up work projects gradients with random matrices [22] or low-rank approximations [25] at the cost of reduced accuracy. Representation-based methods output fixed-size embeddings, which are more manageable in size.
@@ Appendix E Details of Baselines
LoGra is a toolkit for efficient influence function calculation (https://github.com/logix-project/logix), and its core techniques are introduced in Choe et al. [16]. Its key idea is to use low-rank gradient projection to achieve implicit gradient projection during backpropagation, thereby reducing computational cost. We experimented with different configurations and found that the following variants achieved the best performance in our evaluation (as also described in Eq. 3): (i) applying the LoGra module to all MLP layers, (ii) initializing the LoGra weights using PCA, (iii) approximating the Hessian using KFAC, and (iv) performing unit normalization of vectors.
@@
TracIn [22] is derived from a first-order approximation of model weights and essentially computes the dot product between gradients without Hessian correction. For implementation efficiency, we use LoGra to perform gradient projection and apply unit normalization to the vectors. The original TracIn keeps multiple checkpoints and ensembles their influence scores, but this approach linearly increases computation and storage costs. Instead, we use only one checkpoint (i.e., the final checkpoint) of the warmed-up LM.
@@
LESS [25] adopts LoRA tuning instead of full fine-tuning and projects the LoRA gradient to a lower dimension using a random matrix. Its key idea is to use AdamW states (first- and second-order momentum) to correct the gradients. We use the official implementation (https://github.com/princeton-nlp/LESS) and set the projection dimensions to 768 and 8192, which are the default dimensions in the LESS paper.
@@
TRAK [12] is a representative gradient-based method that provides an efficient implementation of the influence function. TRAK approximates the Hessian using the Fisher Information Matrix (FIM), ensembles influence scores across multiple checkpoints, and adjusts the influence score of each data sample by multiplying it with the model’s prediction probability. DsDm [33] later applied TRAK to LLM data selection tasks. Since DsDm has not released its official code (https://github.com/MadryLab/DsDm), we approximate its implementation (i.e., TRAK) using the Logix framework, where the Hessian is approximated with FIM (also known as raw Hessian approximation in Logix), LoGra is used to project the gradient to different dimensions with random weight initialization, and only the final checkpoint is applied without checkpoint ensembling.
@@
RDS was introduced in Hanawa et al. [6], which uses the model’s hidden states as data representations. Through an evaluation of different representation design choices, we use average pooling over the last-layer hidden states of all tokens as the sentence embedding.
@@
GTE [35] is a general text embedding model trained on a mixture of text retrieval, text classification, text clustering, and text entailment data. It achieves strong performance on text embedding leaderboards relative to its model size (e.g., https://huggingface.co/spaces/mteb/leaderboard). For direct evaluation, we assessed various GTE model sizes, including Small, Base, Large, 1.5B, and 7B.
@@
AirRep is inspired by Datamodels [1], in terms of problem formulation for data attribution and both methods focus on estimating empirical group influence scores. A key difference lies in how embeddings are generated. Datamodels assign fixed embedding vectors for all instances in a given data collection, and require the full retraining when new instances are added. AirRep, on the other hand, uses a trained encoder to produce the embeddings for any new instance on the fly. As a result, AirRep is computationally much more efficient than Datamodels in handling dynamically changing data. Empirical findings in [12] shows that Datamodels typically incur inference costs around 100x greater than gradient-based methods to achieve comparable performance. AirRep addresses this issue with the trained encoder, and being 80x more efficient when handling new instances. We will add these discussions in our revised version of the paper.

### v2

@@ 5 Evaluation of Data Attribution
To evaluate these data attribution methods, we follow [Ilyas2022DatamodelsPP] and use the linear datamodeling score (LDS) as our evaluation metric. For each evaluation dataset, we sample 100 different random subsets $(S_{1},\cdots,S_{100})$ of its training set, each containing $n$ training examples, and train the target LM on each of these subsets. For each example $x$ in the test set, we approximate the expectation of the model prediction $r(x,S_{i})$ as the evaluation cross-entropy loss. Given the score $f(x,S_{i})$ predicted by the data attribution model, the LDS is computed as the Spearman rank correlation between the true and estimated influence scores: $\rho\bigl(\{r(x,S_{i}):i\in[100]\},\{f(x,S_{i}):i\in[100]\}\bigr)$. Finally, the LDS on the dataset is averaged across all test examples.
@@ Results
[…]
@@ 7 Evaluation of Data Classification
TDA methods have been applied to measuring data similarity for the purpose of model explainability [Hanawa2021EvaluationOS]. To evaluate the effectiveness of such explanations, Hanawa et al. [Hanawa2021EvaluationOS] introduced the “Identical Class Test” for image classification, which assesses whether the most influential training example identified by TDA methods belongs to the same class as the test instance.
@@ Results
Table 2 presents the evaluation results of LoGra, GTE-Small, and AirRep. [^1] Here we focus on a subset of most competitive or directly relevant baselines, see Appendix H.2 for more. The results indicate that AirRep significantly improves data classification accuracy on FLAN compared to GTE-Small (increasing from 50.59 to 86.41) and also outperforms LoGra. Notably, since AirRep is trained without using any data labels, this suggests that the model can learn to represent task-related information in an unsupervised manner. A similar trend is observed for Tulu, where AirRep demonstrates superior performance in identifying the data source. For SafeRLHF, AirRep outperforms LoGra but lags behind GTE-Small, likely because AirRep’s training data, generated from UltraChat, does not contain harmful content and thus lacks the necessary learning signal.
@@ 9 Amortizing Training Cost
Compared with gradient-based methods, AirRep is more efficient at inference but requires training on auto-generated data, which incurs additional cost. Our evaluation above shows that FLAN- or UltraChat-pretrained AirRep generalizes well on new TDA data and tasks. We now consider another setting where AirRep is re-trained from scratch for each TDA task, and analyze how its inference efficiency can amortize its training cost [Covert2024StochasticAA].
@@
Then, Figure 6 (right) compares the number of training examples that can be processed under different A100 GPU hour budgets by AirRep (crossover checkpoint) and LoGra, accounting for both data generation and model training costs. LoGra completes model training earlier, but AirRep—despite longer training time—achieves nearly two orders of magnitude more throughput once model is trained. Notably, a turning point appears around 475K examples, where both methods take 5.6 GPU hours. Beyond this, even with retraining, AirRep is faster than LoGra. For example, with 24 GPU hours, AirRep can attribute over 100M examples, whereas LoGra can process only around 2M. This highlights the advantage of AirRep for large-scale TDA scenarios. We additionally emphasize that, as shown in Section˜5, AirRep can be transferred from a small model to larger models and still attain stronger performance than all baselines, so the “re-training from scratch” results in this section represent worst-case results for AirRep.
@@
Figure 7: Evaluate with different target LMs. ArRep shows strong generalization.
Amortizing Across Models
We evaluate AirRep, trained with Qwen-0.5B-generated supervision, using different target LMs (Qwen-0.5B, Qwen-7B, Llama-1B, Qwen3-0.6B, Qwen3-0.6B (Thinking), TinyLlama-1B, GPT-2) on FLAN. Results in Figure 7 show that across all targets, AirRep consistently surpasses LoGra, even when LoGra computes gradients on the target LMs. We also observe that LoGra is sensitive to the proxy used for gradient calculation—when the target model differs significantly (e.g., GPT-2), its performance drops notably, whereas AirRep maintains strong performance. This indicates the robust generalization of AirRep and that its training cost can be amortized across different target language models.
@@ 10 Related Work
Most existing work in TDA focuses on gradient-based methods, such as Influence Functions [Koh2017UnderstandingBP] and their subsequent optimizations, including group influence estimation [Koh2019OnTA, Basu2019OnSG], Hessian approximation [Schioppa2021ScalingUI, Grosse2023StudyingLL, Kwon2023DataInfEE], gradient normalization [Barshan2020RelatIFIE, Xia2024LESSSI], gradient projection [Pruthi2020EstimatingTD, Choe2024WhatIY], model ensembling [Park2023TRAKAM, Engstrom2024DsDmMD, Wang2024CapturingTT], and distillation of the proxy influence models [Ladhak2022ContrastiveEA, Liu2024OnTD, Yu2024MATESMD]. However, studies have found that the derivation of gradient-based methods may rely on incorrect assumptions in large neural networks [Bae2022IfIF, Basu2020InfluenceFI]. Additionally, gradient-based methods suffer from the computational expense of gradient calculations and Hessian inverses [Grosse2023StudyingLL, Engstrom2024DsDmMD].
Alternatively, representation-based approaches for TDA measure data influence in the representation space [Zhang2018TheUE, Rajani2020ExplainingAI, Hanawa2021EvaluationOS, Akyrek2022TowardsTF], offering computational and storage efficiency. However, current methods rely on heuristically designed representations without task-aware and model-specific optimization, often leading to suboptimal results [Park2023TRAKAM]. Some research on TDA explores simulation-based methods such as datamodels  [Feldman2020WhatNN, Ghorbani2019DataSE, Jia2019TowardsED, Ilyas2022DatamodelsPP], which are usually not tractable for LLM applications due to their computational expense.
TDA methods have been applied to various purposes in machine learning, such as identifying mislabeled training data points [Koh2017UnderstandingBP, Kong2022ResolvingTB], detecting data poisoning attacks [Fang2020InfluenceFB, Jagielski2020SubpopulationDP], and guiding data augmentation [Lee2020LearningAN, Oh2021InfluenceguidedDA]. TDA methods have also been used for understanding LLM generalization [Grosse2023StudyingLL] and learning processes [Chang2024ScalableIA, Ruis2024ProceduralKI]. Another notable line of related work is data selection for LLMs, which focuses on curating data to optimize test performance [Albalak2024ASO]. TDA-like functions for measuring relationships between data have been widely adopted in data selection pipelines to enhance the quality and diversity of the selected data. Both gradient-based [Thakkar2023SelfInfluenceGD, Xia2024LESSSI, Yu2024MATESMD] and representation-based [Xie2023DataSF, Parkar2024SelectLLMCL, Shao2024BalancedDS, Das2023DEFTUCSDE] TDA approaches have been applied. Meanwhile, studies have explored using text embedding models to select in-context examples for LLM prompting [Li2023UnifiedDR, Wang2023LearningTR, Rubin2021LearningTR].
@@ 11 Conclusion
This paper introduces a new approach to training data attribution called AirRep, which combines the advantages of task-driven optimization in gradient-based approaches with the efficiency of representation-based methods. AirRep represents training data as embeddings and scores their influence using inner-product similarity in the representation space. AirRep’s novel attention-based pooling mechanism captures the group effects, and the entire model is optimized using a novel proposed weighted pairwise ranking objective over automatically generated influence signals. Evaluation on LLM fine-tuning demonstrates the effectiveness and generalization of AirRep.
Limitations of this work include the added cost of model training and evaluation limited to the LLM fine-tuning stage. While our current evaluation focuses on language tasks, the design of AirRep is modality-agnostic—it relies only on the availability of suitable encoders and the ability to estimate loss differences. In principle, this framework could be applied to other domains such as vision or multimodal settings. For future work, we plan to extend our method to additional scenarios, such as the pre-training stages of LLMs and multi-modal model training.
@@ B.1 Group Influence Function
Extended from Basu et al. [Basu2019OnSG] results, we show that the structure of the high-order group influence function is given by:
@@ B.2 Deriving the First-Order Group Influence Function
$0\;\approx\;0\;+\;\left(\!\frac{1}{n-m}\,\sum_{z_{i}\in S}\nabla_{\theta}\ell(z_{i},\theta^{*})\right)\epsilon\;+\;\mathbf{H}\,\Delta_{\epsilon}.$
@@
$f_{\text{IF}}(x,S)=\nabla_{\theta}\ell\bigl(x,\theta^{*}\bigr)^{\top}\mathbf{H}^{-1}\;\;\sum_{z_{i}\in S}\;\;\nabla_{\theta}\ell(z_{i},\theta^{*})$
@@ B.4 Deriving the Third-Order Group Influence Function
For many practical purposes, second‐order corrections already capture the leading “beyond first‐order” effects. However, in order to have a more complete perception of the structure of the influence function (and also to address our curiosity), we will continue to derive the third-order group influence function.
@@
    $\text{(i)}=\left.\frac{\partial^{3}}{\partial\epsilon^{3}}\nabla_{\theta}L_{S}(\theta^{*})\right|_{\epsilon=0}=0.$
@@
    $\text{(ii)}=\nabla^{2}_{\theta}L_{S}(\theta^{*},0)\,\Delta_{3}=\mathbf{H}\,\Delta_{3}.$
  • Term (iii) vanishes because the per‐sample weight is linear in $\epsilon$:
    $\text{(iii)}=\frac{1}{2}\left.\frac{\partial^{2}}{\partial\epsilon^{2}}\nabla^{2}_{\theta}L_{S}(\theta^{*},0)\right|_{\epsilon=0}\,\Delta_{1}=0.$
@@
    $\text{(iv)}=\left.\frac{\partial}{\partial\epsilon}\nabla^{2}_{\theta}L_{S}(\theta^{*},0)\right|_{\epsilon=0}\,\Delta_{2}.$
@@
  • Term (v) involves the third derivative with respect to $\theta$. In most influence‐function derivations, one often assumes that higher‐order derivatives $\nabla^{3}_{\theta}L(\theta^{*})$ are negligible. Following this convention, we consider:
    $\text{(v)}=\Delta_{1}^{\top}\,\nabla^{3}_{\theta}L_{S}(\theta^{*},0)\,\Delta_{2}=0.$
@@
    $\text{(vi)}=\frac{1}{2}\,\Delta_{1}^{\top}\,\left.\frac{\partial}{\partial\epsilon}\nabla^{3}_{\theta}L_{S}(\theta^{*},0)\right|_{\epsilon=0}\,\Delta_{1}=0.$
@@ B.5 Deriving the Higher-Order Group Influence Function
We see that at each order $k$ in $\epsilon$, the only “new” term that enters the equations (and thus the recursion for $\Delta_{k}$) is the first derivative with respect to $\epsilon$ multiplied by lower‐order $\Delta_{j}$’s, plus the usual expansions in powers of $\Delta_{\epsilon}$ (where $\Delta_{\epsilon}=\theta_{\epsilon,S}^{*}-\theta^{*}$).
@@ Appendix C Comparison between gradient-based and representation-based methods
  • Computation cost: Gradient-based methods generally rely on the computation of gradients and some approximation of the Hessian, making them computationally expensive. In practice, the cost of gradient-based methods is comparable to training the model on the all of the target examples because computing influence involves computing loss gradients for each target example. In contrast, representation-based methods are efficient and applicable for large-scale retrieval [Johnson2017BillionScaleSS].
  • Storage cost: Influence functions can be particularly storage-intensive as they can involve storing the full gradient (equal in size to the model parameters themselves) for each data sample, which prevents large-scale data attribution. Follow-up work projects gradients with random matrices [Pruthi2020EstimatingTD] or low-rank approximations [Xia2024LESSSI] at the cost of reduced accuracy. Representation-based methods output fixed-size embeddings, which are more manageable in size.
@@ Appendix E Details of Baselines
LoGra is a toolkit for efficient influence function calculation (https://github.com/logix-project/logix), and its core techniques are introduced in Choe2024WhatIY. Its key idea is to use low-rank gradient projection to achieve implicit gradient projection during backpropagation, thereby reducing computational cost. We experimented with different configurations and found that the following variants achieved the best performance in our evaluation (as also described in Eq. 3): (i) applying the LoGra module to all MLP layers, (ii) initializing the LoGra weights using PCA, (iii) approximating the Hessian using KFAC, and (iv) performing unit normalization of vectors.
@@
TracIn [Pruthi2020EstimatingTD] is derived from a first-order approximation of model weights and essentially computes the dot product between gradients without Hessian correction. For implementation efficiency, we use LoGra to perform gradient projection and apply unit normalization to the vectors. The original TracIn keeps multiple checkpoints and ensembles their influence scores, but this approach linearly increases computation and storage costs. Instead, we use only one checkpoint (i.e., the final checkpoint) of the warmed-up LM.
@@
LESS [Xia2024LESSSI] adopts LoRA tuning instead of full fine-tuning and projects the LoRA gradient to a lower dimension using a random matrix. Its key idea is to use AdamW states (first- and second-order momentum) to correct the gradients. We use the official implementation (https://github.com/princeton-nlp/LESS) and set the projection dimensions to 768 and 8192, which are the default dimensions in the LESS paper.
@@
TRAK [Park2023TRAKAM] is a representative gradient-based method that provides an efficient implementation of the influence function. TRAK approximates the Hessian using the Fisher Information Matrix (FIM), ensembles influence scores across multiple checkpoints, and adjusts the influence score of each data sample by multiplying it with the model’s prediction probability. DsDm [Engstrom2024DsDmMD] later applied TRAK to LLM data selection tasks. Since DsDm has not released its official code (https://github.com/MadryLab/DsDm), we approximate its implementation (i.e., TRAK) using the Logix framework, where the Hessian is approximated with FIM (also known as raw Hessian approximation in Logix), LoGra is used to project the gradient to different dimensions with random weight initialization, and only the final checkpoint is applied without checkpoint ensembling.
@@
RDS was introduced in Hanawa2021EvaluationOS, which uses the model’s hidden states as data representations. Through an evaluation of different representation design choices, we use average pooling over the last-layer hidden states of all tokens as the sentence embedding.
@@
GTE [Li2023TowardsGT] is a general text embedding model trained on a mixture of text retrieval, text classification, text clustering, and text entailment data. It achieves strong performance on text embedding leaderboards relative to its model size (e.g., https://huggingface.co/spaces/mteb/leaderboard). For direct evaluation, we assessed various GTE model sizes, including Small, Base, Large, 1.5B, and 7B.
@@
AirRep is inspired by Datamodels [Ilyas2022DatamodelsPP], in terms of problem formulation for data attribution and both methods focus on estimating empirical group influence scores. A key difference lies in how embeddings are generated. Datamodels assign fixed embedding vectors for all instances in a given data collection, and require the full retraining when new instances are added. AirRep, on the other hand, uses a trained encoder to produce the embeddings for any new instance on the fly. As a result, AirRep is computationally much more efficient than Datamodels in handling dynamically changing data. Empirical findings in [Park2023TRAKAM] shows that Datamodels typically incur inference costs around 100x greater than gradient-based methods to achieve comparable performance. AirRep addresses this issue with the trained encoder, and being 80x more efficient when handling new instances. We will add these discussions in our revised version of the paper.